Find the equation of the line joining the points (0, 5) and (2, 11).
Thank's for your help.

Found 3 solutions by stanbon, CharlesG2, edjones:
Answer by stanbon(75887)

Find the equation of the line joining the points (0, 5) and (2, 11).
---
Form: y = mx + b
You need to determine m and b
---
But "b" is the y- intercept and (0,5) is the y-intercept, so b = 5
---
m = (11-5)/(2-0) = 3
---
So the equation is y = 3x + 5
=========================
Cheers,
Stan H.

Answer by CharlesG2(834)

Find the equation of the line joining the points (0, 5) and (2, 11).
Thank's for your help.
slope m = y2-y1 / x2-x1
m = (11-5)/(2-0)
m = 6/2 = 3
a line is y=mx+b , where m is the slope and b is the y-intercept (0,b)
y=3x+b solve for b
y=3 (0)+b
y=0+b
y=b that is when b=5
y=3x+5

Answer by edjones(8007)

(11-5)/(2-0)=6/2=3=m the slope
y-2=3(x-0)
y=3x+2
.
Ed